Assume that a neutron breaks into a proton and an electron. The energy released during this process is (Mass of neutron = 1.6725 x 10–27kg; mass of proton = 1.6725 x 10–27kg; mass of electron = 9 x 10–31 kg)
A. 7.10 Mev
B.6.30 Mev
C.5.4 Mev
D.0.73 Mev

Assume that a neutron breaks into a proton and an electron. The energy released during this process is (Mass of neutron = 1.6725 x 10–27kg; mass of proton = 1.6725 x 10–27kg; mass of electron = 9 x 10–31 kg)

Grade:

1 Answers

Arun
5 years ago
Dear student
 
Delta m = mp + me – mn
 
= 9 * 10^-31 kg
 
energy released =  9 * 10^-31 kg * c^2
 
when converting it in Mev = 0.73 Mev